b. Find the series solution of the following differential equation


dz/dt -zet=tz


1
Expert's answer
2021-06-23T06:56:36-0400

Solution

Let

"z\left(t\right)=\sum_{n=0}^{\infty}{a_nt^n}"

"\frac{dz}{dt}=\sum_{n=1}^{\infty}{na_nt^{n-1}}=\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}t^n}"

It’s known that

"e^t=\sum_{n=0}^{\infty}\frac{t^n}{n!}"

"ze^t=\sum_{n=0}^{\infty}\left(\sum_{i=0}^{n}\frac{a_{n-i}}{i!}\right)t^n"

Plug these into the differential equation

"\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}t^n}-\sum_{n=0}^{\infty}\left(\sum_{i=0}^{n}\frac{a_{n-i}}{i!}\right)t^n=\sum_{n=1}^{\infty}{a_{n-1}t^n}"

Coefficient near tn are equal to zero.

n=0 : a1- a0=0 => a1 = a0 

n=1 : 2a2- a0- a1 = a0 => a2= a0+ a1/2 => a2= 3a0/2    

n=2 : 3a3- a0/2- a1 -a2= a1 => a3= a0/3!+2a1/3+ a2/3 => a3= a0/3!+2a0/3+ 3a0/6 => a3= 8a0/3!     

n=3 : 4a4- a0/3!- a1 /2!-a2-a3 =a2 => a4= a0/4!+a1/8+ 2a2/4 + a3/4  => a4= a0/4!+a0/8+ 2*3a0/8 + 8a0/4!  => a4= 30a0/4!   

n=4 : 5a5- a0/4!- a1 /3!-a2/2!-a3 -a4 =a3 => a5= a0/5!+ a1 /(5*3!)+a2/(5*2!)+2a3/5+a4/5 => a5= a0/5!+ a0/(5*3!)+3a0/(5*2*2!)+16a0/(5*3!)+30a0/5! = a0/5!+ 4a0/5!+18a0/5!+64a0/(5!)+30a0/5! => a5= 117a0/5!  

And so on. For an= bn/n!  we’ll get simplified recurrent formula for bn  

"b_1=b_0;\ b_{n+1}=\sum_{i=0}^{n}{\binom{n}{i}b_{n-i}}+nb_{n-1}\ (n>0)"

and solution

"z\left(t\right)=\sum_{n=0}^{\infty}{\frac{b_n}{n!}t^n}"